# More Women, Same Stereotypes: Unpacking the Gender Bias Paradox in Large Language Models

### v2

@@ 1 Introduction
To address these shortcomings, we propose a simple evaluation strategy that avoids predefined scenarios and conceals the intent of the test. Specifically, we request LLMs to generate a story opening and setting for a character in a specific occupation and analyze the gender cues in the output. We execute the exact prompt multiple times for each occupation and evaluate the gender distribution across repeated outputs. Thus, we eliminate the need for strictly designed scenarios or carefully tailored prompts, allowing biases to appear naturally within free-form storytelling, and avoid explicitly directing the model’s attention toward gender-related considerations, providing an authentic view of occupational gender stereotypes.
Based on this new evaluation strategy, we have the following observations. First, LLMs exhibit an apparent overrepresentation of female characters in their generated stories, with 35 of 106 occupations showing female dominance ($80\%$ or more characters are female) across the 10 LLMs tested, whereas male characters did so in only 5 out of 106. Interestingly, this female dominance is not limited to traditionally stereotypical female-associated occupations (Section 4.1). A possible reason for this pattern is that alignment techniques, such as supervised fine-tuning (SFT) and reinforcement learning from human feedback (RLHF) employed by model developers, may have influenced the models towards increased female representation. This effort could lead to the overrepresentation of female characters in a wide range of occupations, not only those traditionally associated with women. We validated this assumption by comparing the models with and without SFT and RLHF (Section 4.3). Second, the rankings of occupations by gender ratios in LLM outputs align more closely with the perceived gender associations of occupations (as measured by human ratings) than the actual gender distributions in the labor force. This indicates that LLMs are more likely to mirror societal perceptions of gender roles than real-world demographic distributions, potentially amplifying biases rooted in societal stereotypes (Section 4.2). Although this observation may seem contradictory to the first conclusion, it is not. The increased representation of females across all categories appears to lead many occupations to become female-dominated. However, when occupations are ranked by gender composition ratios, the resulting sequence still strongly aligns with human perceptions of gender associations.
@@ 2.1 Evaluating Bias by LLMs’ Decisions on Strictly Designed Scenarios
2.2 Evaluating Bias by LLMs’ Word Choices on Carefully Designed Prompts
@@ 3 Evaluating Occupational Stereotypes
We selected Llama 3.2 3B (Meta AI, 2024), Gemma 2 (2B, 9B and 27B) (Google et al., 2024a), GPT-4o, GPT-4o mini (OpenAI et al., 2024), Gemini 1.5 Flash and Gemini 1.5 Flash-8B, Gemini 1.5 pro and Gemini 2.0 flash (Google et al., 2024b) for evaluation because these models strike a balance between accessibility and representational capabilities, allowing us to conduct comprehensive evaluations without requiring complex infrastructure. Specifically, we built Llama 3.2 3B and Gemma 2 (2B, 9B, and 27B) via Ollama, while the GPT-4o, GPT-4o mini, Gemini 1.5 Flash, Gemini 1.5 Flash-8B, Gemini 1.5 Pro, and Gemini 2.0 Flash models were accessed through official APIs. All models were evaluated using default settings, as this is the configuration that most users experience.
@@ 3.4 Bias Analysis
To evaluate the gender stereotypes exhibited by each LLM, we compared the gender distribution of the story characters across occupations against two benchmarks: (1) the stereotype gender associations (Kennison and Trofe, 2003), which reflect human perceptions of gender stereotypes for occupations; (2) the gender distribution in the U.S. labor force (U.S. Bureau of Labor Statistics, 2023), representing real-world demographics. These comparisons assess whether LLMs’ gender biases align with societal perceptions, real-world statistics, or diverge in ways that introduce new biases.
@@
Table 1: Kendall’s Tau and Averaged Male Proportion Deviations between LLMs and two benchmarks (U.S. BLS and GSR). For Kendall’s Tau, higher is more consistent. For Averaged Male Proportion Deviations, lower means a bias towards females. All Kendall’s Tau results are significant ($p$-values $\ll 10^{-5}$).
@@ 4 Discovery
Our experiments revealed notable and complex patterns in LLMs’ behaviors regarding gender representation in occupational narratives.
@@ 4.1 Pervasive Overrepresentation of Females
Figure 1 ranks occupations based on human ratings, from those perceived as the most female-oriented to those considered the most male-oriented based on GSR. The heatmap shows that all LLMs tested show a significant tendency toward female characters. This pronounced trend has direct consequences at the occupational level: of the 106 occupations analyzed, 35 featured female characters in at least $80\%$ of the generated stories, while only 5 occupations exhibited similar dominance for male characters. Figure 2 further illustrates this gender distribution across all occupations. Although the median gender ratios of real-world data (US BLS) and human evaluations (GSR) are relatively balanced (0.473 and 0.565 male proportion, respectively), all LLMs exhibit a significant bias, with median male proportions falling below $20\%$, and in some cases even below $10\%$.
@@ 4.3 Female Overrepresentation May Stem from Alignment
This discrepancy implies that strong female overrepresentation may be an emergent property or an intentional outcome after alignment. Developers, in their efforts to mitigate pre-existing biases or to achieve certain diversity goals, might have calibrated the models to over-generate female characters. Although such interventions may aim for increased inclusivity, they risk creating new artificial skews.

### v3

@@ 1 Introduction
To address these shortcomings, we propose a simple evaluation strategy that avoids predefined scenarios and conceals the intent of the test. Specifically, we request LLMs to generate a story opening and setting for a character in a specific occupation and analyze the gender cues in the output. We execute the exact prompt multiple times for each occupation and evaluate the gender distribution across repeated outputs. Thus, we eliminate the need for strictly designed scenarios or carefully tailored prompts, allowing biases to appear naturally within free-form storytelling and avoiding directing the model’s attention toward gender-related considerations, providing an authentic view of occupational gender stereotypes.
Based on this new evaluation strategy, we observe the following. First, LLMs exhibit an overrepresentation of female characters in their generated stories, with 35 of 106 occupations showing female dominance ($80\%$ or more characters are female) across the 10 LLMs tested, whereas male characters did so in only 5 out of 106. Interestingly, this female dominance is not limited to traditionally stereotypical female-associated occupations (Section 4.1). A possible reason for this pattern is that alignment techniques, such as supervised fine-tuning (SFT) and reinforcement learning from human feedback (RLHF) employed by model developers, may have influenced the models towards increased female representation. This effort could lead to the overrepresentation of female characters in a wide range of occupations, not only those traditionally associated with women. We validated this assumption by comparing the models with and without SFT and RLHF (Section 4.3). Second, the rankings of occupations by gender ratios in LLM outputs align more closely with the perceived gender associations of occupations (as measured by human ratings) than the actual gender distributions in the labor force. This indicates that LLMs are more likely to mirror societal perceptions of gender roles than real-world demographic distributions, potentially amplifying biases rooted in societal stereotypes (Section 4.2). Although this observation may seem contradictory to the first conclusion, it is not. The increased representation of females across all categories appears to lead many occupations to become female-dominated. However, when occupations are ranked by gender composition ratios, the resulting sequence still strongly aligns with human perceptions of gender associations.
@@ 2.1 Evaluating Bias by LLMs’ Decisions on Strictly Designed Scenarios
2.2 Evaluating Bias by LLMs’ Word Choices
@@ 3 Evaluating Occupational Stereotypes
We selected Llama 3.2 3B (Meta AI, 2024), Gemma 2 (2B, 9B and 27B) (Team et al., 2024a), GPT-4o, GPT-4o mini (OpenAI et al., 2024), Gemini 1.5 Flash and Gemini 1.5 Flash-8B, Gemini 1.5 pro and Gemini 2.0 flash (Team et al., 2024b) for evaluation because these models strike a balance between accessibility and representational capabilities, allowing us to conduct comprehensive evaluations without requiring complex infrastructure. Specifically, we built Llama 3.2 3B and Gemma 2 (2B, 9B, and 27B) via Ollama, while the GPT-4o, GPT-4o mini, Gemini 1.5 Flash, Gemini 1.5 Flash-8B, Gemini 1.5 Pro, and Gemini 2.0 Flash models were accessed through official APIs. All models were evaluated using default settings, as this is the configuration that most users experience.
@@ 3.4 Bias Analysis
To evaluate the gender stereotypes exhibited by each LLM, we compared the gender distribution against two benchmarks: (1) the stereotype gender associations (Kennison and Trofe, 2003), reflecting human perceptions of gender stereotypes for occupations; (2) the gender distribution in the U.S. labor force (U.S. Bureau of Labor Statistics, 2023), representing real-world demographics. These comparisons assess whether LLMs’ gender biases align with societal perceptions, real-world statistics, or diverge in ways that introduce new biases.
@@
Table 1: Kendall’s Tau and Averaged Male Proportion Deviations between LLMs and two benchmarks U.S. BLS and GSR. A higher value of Kendall’s Tau means more consistent; a lower value of Averaged Male Proportion Deviation means a bias towards females. All Kendall’s Tau results are significant ($p$-values $\ll 10^{-5}$).
@@ 4 Discovery
Our experiments revealed notable and complex patterns in LLMs regarding gender representation in occupational narratives.
@@ 4.1 Pervasive Overrepresentation of Females
Figure 1 ranks occupations based on human ratings, from those perceived as the most female-oriented to the most male-oriented based on GSR. The heatmap shows that all LLMs tested show a significant tendency toward female characters. This pronounced trend has direct consequences at the occupational level: of the 106 occupations analyzed, 35 featured female characters in at least $80\%$ of the generated stories, while only 5 occupations exhibited similar dominance for male characters. Figure 2 further illustrates this gender distribution across all occupations. Although the median gender ratios of real-world data (US BLS) and human evaluations (GSR) are relatively balanced (0.473 and 0.565 male proportion, respectively), all LLMs exhibit a significant bias, with median male proportions falling below $20\%$, and in some cases even below $10\%$.
@@ 4.3 Female Overrepresentation May Stem from Alignment
This discrepancy implies that strong female overrepresentation may be an emergent property or an intentional outcome after alignment. Developers, in their efforts to mitigate pre-existing biases or achieve diversity goals, might have calibrated the models to over-generate female characters. Although such interventions may aim for increased inclusivity, they risk creating new artificial skews.